OpenStudy (anonymous):

Can someone solve this please? the points (2,6) and (3,18) lie on the curve y=ax^n use logarithms to find the values of a and n, giving your answers correct to 2d.p. thanks :)

OpenStudy (anonymous):

Substitute the two points in the given equation, gives us the two following equations: \[6=a(2)^n \rightarrow(1)\] and\[18=a(3)^n \rightarrow(2)\] Now, we should solve the two equations for the two unknowns a and n

OpenStudy (anonymous):

Using the first equation we have: \[a={6 \over 2^n}\rightarrow(*)\] Substitute (*) into equation (2): \[18={6 \over 2^n}(3)^n \implies 3=({3 \over 2})^n \implies \log(3)=n(\log({3 \over 2}))\implies n={\log(3) \over \log({3 \over 2})}\]

OpenStudy (anonymous):

That's n=2.71. Now substitute this value of n into (*) to get a.

OpenStudy (anonymous):

So \[a={6 \over 2^n}={6 \over 2^2.71}\approx0.92\]

OpenStudy (anonymous):

where did the 1 and 2 come from?

OpenStudy (anonymous):

how did you get to \[3=(3/2)^n\]